OpenStudy (mathmusician):

The Region R is bounded by the x-axis, x=1, x=3, and y=1/x^3.

OpenStudy (mathmusician):

Find the value of h, such that the vertical line x=h divides the region R into two regions of equal area.

OpenStudy (mathmusician):

@FaiqRaees

OpenStudy (math&ing001):

I guess it's gonna be something like this: \[\int\limits_{1}^{h} \frac{ 1 }{ x ^{3} }dx=\int\limits_{h}^{3} \frac{ 1 }{ x ^{3} }dx\]

OpenStudy (mathmusician):

Thanks

OpenStudy (mathmusician):

May I ask how to solve for h?

OpenStudy (math&ing001):

Solve the integrals, then you'll have an equation of h.

OpenStudy (mathmusician):

OKay I'll try

OpenStudy (math&ing001):

Let me know =)

OpenStudy (mathmusician):

This is as far as I got 

OpenStudy (mathmusician):

@math&ing001

OpenStudy (math&ing001):

Yeah, you're on the right track. Now you just plug in to get the equation.

OpenStudy (mathmusician):

Plug in what?

OpenStudy (math&ing001):

\[[-1/2x^{2}]_{1}^{h}=-\frac{ 1 }{ 2h ^{2} }+\frac{ 1 }{ 2 }\] Same goes for the RHS.

OpenStudy (mathmusician):

\[1 = -\frac{ 1 }{ 2h ^{2} }+\frac{ 1 }{ 2 }\]

OpenStudy (mathmusician):

?

OpenStudy (math&ing001):

Come ooon do I have to solve the whole thing \[[-1/2x ^{2}]_{1}^{h} = [-1/2x ^{2}]_{h}^{3}\]\[=>-\frac{ 1 }{ 2h ^{2} }+\frac{ 1 }{ 2 }=-\frac{ 1 }{ 2*9 }+\frac{ 1 }{ 2h ^{2} }\] Now you multiply both sides by h² to get a second degree equation, you solve for h. One of the solutions should be outside the range [1,3] which you discard. That's it you have h.

OpenStudy (mathmusician):

equals about 1.64

OpenStudy (mathmusician):

Is that right?

OpenStudy (math&ing001):

I got h=sqrt(9/5)=1.34